# Beyond Numeric Awards: In-Context Dueling Bandits with LLM Agents

### v3

@@ 4 Algorithm-Enhanced LLMs for Dueling Bandits
Classic DB algorithms based on the Explore-then-Exploit framework, such as Interleaved Filter 2 (IF2) (Yue et al., 2012), are known to be near-optimal, with matching regret upper and lower bounds up to multiplicative constants. To address the challenges identified in Section 3.2 of using standalone LLM agents for DB, we propose an algorithm-enhanced approach: LLM with Enhanced Algorithmic Dueling (LEAD) to demonstrate the possibility of integrating off-the-shelf DB algorithms with LLM agents through fine-grained adaptive interplay. Our framework, LEAD, enjoys both a regret guarantee and strong empirical performance.
@@ 4.1 Algorithmic Design of LEAD
In this section, we present the design intuitions of LEAD. We begin by discussing the limitations of a naive intervention approach and the desirable properties for an effective Algorithm-Enhanced LLM framework. Based on these considerations, we propose an agentic framework design LEAD, where we can incorporate any Explore-then-Exploit DB algorithms (Zoghi et al., 2014a). As an illustrative example, we use IF2 (Yue et al., 2012) to demonstrate how off-the-shelf algorithms can be integrated within LEAD and provide a detailed description.
@@
As a result, we find that the Explore-Then-Exploit structure is particularly well-suited for LLMs (see Appendix B.1 for a detailed illustration). By selecting an Explore-Then-Exploit DB algorithm as the foundation inside LEAD, we address [Q1]. As an example, we use IF2 as the base to illustrate the theoretical guarantee and empirical performance. This approach can be applied similarly to other algorithms with regret bounds in the Explore-Then-Exploit family.
Algorithmic Framework. The procedures of the LEAD are illustrated in Figure 4 and presented in Algorithm 1 (see more details in Appendix B.2). LEAD (IF2 base) maintains a confidence parameter $\delta$ and a threshold parameter $\epsilon$ that control the algorithm’s confidence of matches between arms. The key components of LEAD (IF2 base) are as follows:
  • Phase 1 (LLM Phase): Utilization of LLM recommended arms: The agentic framework maintains a set of candidate arms $B$. Given two arms suggested by an LLM agent, the framework begins with finding a winner between them, denoted by $b_{\mathrm{LLM}}$. The winning arm $b_{\mathrm{LLM}}$ is then matched with each remaining arm $b\in B$. This phase continues until $b_{\mathrm{LLM}}$ is defeated or all arms in $B$ have been matched. The variable $\mathsf{TrustLLM}$ is used to control the execution of the LLM phase, and it is set to $\mathtt{False}$ when $b_{\mathrm{LLM}}$ is defeated by another arm, indicating that the LLM’s suggestions are no longer trusted.
  • Phase 2 (IF2 Phase): Roll back to IF2: If $b_{\mathrm{LLM}}$ is defeated, the framework switches to implementing one round of IF2 with an incumbent arm $b_{\text{{IF2}}}$ selected based on an estimated preference matrix $\hat{P}$.
@@
4.2 Theoretical Guarantees for LEAD
In this section, we begin by characterizing the vulnerability of using standalone LLM agents for dueling bandits in Theorem 4.1. Then, we provide the theoretical guarantees of LEAD in Theorem 4.2 and  4.3, demonstrating its efficacy and convergence.
Theorem 4.1 (Vulnerability).
The proof of Theorem 4.1 is provided in Appendix B.3.2. The theorem underscores the suboptimality of standalone LLM agents in DB problems, particularly when input prompts are subjected to adversarial attacks. This vulnerability highlights the need for a more robust approach to use in-context LLM agents while offering theoretical guarantees under diverse prompting scenarios.
Regret Bounds. Following the algorithmic design of LEAD in Section 4.1, LEAD (IF2 base) inherits the theoretical guarantees of IF2 (see Appendix B.3.1), while nontrivially leveraging the benefits of LLMs’ exceptional weak regret performance for exploration across a range of instances within RDB. Specifically, LEAD (IF2 base) has the following theoretical guarantee:
Theorem 4.2 (Expected Regret).
$\mathbb{E}\left[\mathsf{WR}(\textsc{LEAD})\right]\leq\min\left\{\widetilde{O}\left(T_{\mathrm{LLM}}+\frac{K\log K}{\epsilon_{1,2}}\right),\widetilde{O}\left(\frac{K\log T}{\epsilon_{1,2}}\right)\right\},$ (3)
where $\widetilde{O}(\cdot)$ hides poly-logarithmic factors of $T$.
The following theorem indicates that the additional term $({K\log K})/{\epsilon_{1,2}}$ in (3) is almost tight. Its proof is provided in Appendix B.3.2.
Theorem 4.3 (Converse).
Given any algorithm ALG for dueling bandits provided with an in-context LLM agent recommending arms, if it satisfies $\mathbb{E}\left[\mathsf{WR}(\textsc{ALG})\right]\leq T_{\mathrm{LLM}}$ for all $T_{\mathrm{LLM}}$, then it must hold ${\mathbb{E}\left[\mathsf{SR}(\textsc{ALG})\right]\geq\mathbb{E}\left[\mathsf{WR}(\textsc{ALG})\right]\geq\Omega\left(T\right)}$ for some instance of the LLM agent.
4.3 Empirical Evaluation of LEAD
4.3.1 Efficacy Evaluation: Strong Regret and Weak Regret
Hyper-parameters. In our implementation of LEAD (see Algorithm 1), there are two hyper-parameters: the threshold parameter $t$, which controls the maximum number of comparisons between arms, and the confidence parameter $\delta$, which determines the confidence level for pruning suboptimal arms. For the threshold parameter $t$, we considered values from the set $\{50,100,200\}$, and for the confidence parameter $\delta$, we explored values from $\{0.1,0.2,0.4\}$. After fine-tuning, we found that setting $t=50$ and $\delta=0.4$ provided the best performance in terms of cumulative regret.
4.3.2 Robustness Evaluation: Noisy and Adversarial Prompts
Recent studies (Loya et al., 2023; Krishnamurthy et al., 2024) have emphasized the importance of varying prompts to elicit the desired behavior from LLMs in decision-making tasks, highlighting the potential limitations of prompt quality. Results obtained from a single prompt template may lead to unreliable conclusions that cannot generalize to real-world situations where optimal prompts are often unavailable. Thus, we evaluate the robustness of LEAD by employing two types of prompt perturbations (see Figure 8) along with the original prompt (see Figure 7). Across all scenarios, LEAD demonstrates superior performance and robustness compared to standalone GPT-4 Turbo.
Figure 6: Duel selection trajectory of GPT-4 Turbo and LEAD under different settings (Figures 7 and 8). Top: original prompt. Middle: noisy prompt (biased history). Bottom: adversarial prompt (reversed goal).
Original Prompt. Under the initial prompt, LEAD leverages the LLM’s ability to quickly identify the best arm through exploration (under the DB instance within RDB Eq.(2)). As shown in Figure 6 (Top Row), we observe that LEAD benefits from the LLM’s exploration ability by initializing with the best arm as the incumbent when entering the IF2 phase. Compared to GPT-4 Turbo, convergence to the Condorcet winner is guaranteed for LEAD with high probability.
Biased History. We inject an incorrect history into the prompt, where each non-optimal arm initially wins against the best arm 10 times, while keeping the underlying preference matrix unchanged. LLM agents are observed to get trapped in local optima for extended periods, where LEAD overcomes this limitation by employing uniform comparisons in the IF2 phase to escape such suboptimal exploration modes.
Reversed Goal. When the prompt is adversarially modified from maximizing reward to minimizing, the LLM consistently recommends non-optimal arms after its exploration stage. Even with adversarial prompts, LEAD still achieves near-optimal cumulative strong regret. Since the LLM’s exploration capability is only utilized within the bounded length of the Match Arms procedure, the impact of the reversed goal on the exploitation phase is mitigated.
Figure 5 (right) presents the cumulative strong regret results comparing LEAD against standalone LLM agents and the IF2 algorithm across three prompt designs. Notably, LEAD with $\delta=1/(TK^{2})$ (consistent with IF2 to showcase the robust behavior) achieves near-optimal cumulative regret with low variance even with noisy and adversarial prompts, validating the regret bounds stated in Theorem 4.2. LEAD and IF2 converge to the best arm within 2000 steps, while GPT-4 Turbo’s cumulative expected regret continues to increase, indicating the instability of standalone in-context LLM agents.
5 Conclusion
This paper evaluates LLMs as in-context decision-makers for standard context-free dueling bandits (DB) with a Condorcet Winner, offering the first systematic insights into their strengths and limitations. Our findings reveal that LLMs’ decision-making in DB, driven by linguistic priors, achieves exceptional weak regret performance across both transitive and intransitive environment instances in the short-term. However, LLMs lack the necessary criteria for convergence and long-term performance generalization to hard scenarios, leading to an optimality gap between LLMs and classic DB algorithms in terms of strong regret. To bridge this gap, we propose LEAD, an agentic flow framework that integrates off-the-shelf DB algorithms with LLM agents through fine-grained adaptive interplay. This framework provides theoretical guarantees and demonstrates robust performance even under noisy and adversarial prompts. Our work contributes to the problem of In-Context Reinforcement Learning (ICRL). The framework we proposed sheds light on how language-based reasoning can inspire robust frameworks that translate words into actions, paving the way for more trustworthy AI systems through the interplay between rule-based experts and in-context LLM agents.
Appendix
  1. A.
    Related Works
  2. B.
    Theoretical Part: Algorithm Design and Analysis of LEAD
  3. C.
    Experimental Part: Prompt Design and Supplementary Results
Appendix A Related Works
We provide the detailed related works as follows.
Dueling Bandits. The problem of dueling bandits was initially introduced in [Yue et al., 2012]. Various methods have been proposed to tackle the task since then. These methods can be broadly classified into two categories as Explore-Then-Exploit methods and Ongoing Regret Minimization methods according to [Zoghi et al., 2014a]. Explore-Then-Exploit methods focus on identifying the best arm with high confidence before exploiting it, such as Interleaved Filter (IF[Yue et al., 2012] and Beat the Mean (BTM[Yue and Joachims, 2011], etc. In contrast, Ongoing Regret Minimization methods explicitly target the objective of minimizing cumulative regret, including Relative Upper Confidence Bound (RUCB[Zoghi et al., 2014b] and Self-Sparring [Sui et al., 2017], etc. Dueling bandit problem and preference feedback in general has a wide variety of applications, including recommendation systems [Yue et al., 2012], robotics [Tucker et al., 2020], and most recently, the training algorithm of large language models, such as Reinforcement Learning from Human Feedback (RLHF) [Ouyang et al., 2022].
LLM Agents for Multi-Armed Bandits. Several recent works have explored evaluating the capabilities of LLMs in bandit problems. For example, [Baheri and Alm, 2023] proposed an approach to enhance contextual bandits by integrating LLMs as encoders. The LLMs’ ability to capture rich semantic and syntactic information from textual contexts is leveraged to provide the algorithm with a more informative representation of the context. The LLM-augmented algorithm transforms the raw context into a latent space vector using the LLM’s encoding capabilities. This encoded context is then used to guide the decision-making process. [Krishnamurthy et al., 2024] investigates whether LLMs can engage in exploration in simple MAB environments without additional training. They compared various prompt designs and found that GPT-4 with zero-shot chain-of-thought (CoT) reasoning and an externally summarized interaction history performed the best, while other configurations failed in exploration, either by never selecting the best arm after initial rounds or by selecting all arms nearly equally often. Different from the previous results, in this work we go beyond the settings of numeric rewards and investigate the capabilities of LLMs under preference feedback.
In-Context LLMs for Decision-Making. Beyond bandit problems, LLM agents have demonstrated strong capabilities in complex reasoning across a wide range of in-context reinforcement learning and decision-making tasks [Laskin et al., 2022, Lee et al., 2024, Zhou et al., 2022, Yao et al., 2024]. Various existing works aim to understand LLM agents’ capabilities for in-context decision-making, with notable examples including planning [Huang et al., 2022, Hao et al., 2023]. Additionally, LLM agents have been shown to enhance embodied agents in robotic applications by providing advanced task planning abilities [Brohan et al., 2023] and reward designing [Ma et al., 2023], further enabling the development of lifelong learning agents [Wang et al., 2023]. Besides these empirical successes, the authors of [Park et al., 2024] analyzed LLMs’ interactions in online learning and game theory settings through the lens of the regret metrics. They identified simple cases where LLMs fail to be no-regret. Another line of research incorporates LLMs into classic decision-making frameworks to create LLM-augmented online decision-makers. For instance, Liu et al. [Liu et al., 2024] utilized LLMs to enhance the components of warm starting, sampling candidates, and surrogate modeling in Bayesian optimization. Our work contributes to this broad area by integrating LLM agents with the classic Explore-then-Exploit DB algorithms to enhance the utilization of preference feedback.
Appendix B Algorithm Design and Analysis of LEAD
B.1 Algorithm Design Logic
Explore-then-Exploit algorithms as Ideal Candidates. Classic DB algorithms can be classified into two categories: Explore-Then-Exploit methods and Ongoing Regret Minimization methods [Zoghi et al., 2014a]. Among these, Explore-Then-Exploit structure stands out as particularly well-suited for LLM augmentation:
  • Its symbolic representation of the algorithm’s logic enables clear integration of LLM suggestions at specific points without disrupting the overall structure and theoretical guarantees. In contrast, algorithms like Self-Sparring in [Sui et al., 2017] are less symbolic, making them less suitable for direct LLM augmentation.
B.2 Detailed Procedure Description
@@
In Procedure 1 below, we describe the Match Arms procedure used in LEAD (see Algorithm 1 and Figure 4).
@@
We reprise the IF2 procedure in [Yue et al., 2012] below to complement the presentation of LEAD.

### v4

@@ 4 Algorithm-Enhanced LLMs for Dueling Bandits
Classic DB algorithms based on the Explore-then-Exploit framework, such as Interleaved Filter 2 (IF2) (Yue et al., 2012), are known to be near-optimal, with matching regret upper and lower bounds up to multiplicative constants. To address the challenges identified in Section 3.1 of using standalone LLM agents for DB, we propose an algorithm-enhanced approach: LLM with Enhanced Algorithmic Dueling (LEAD) to demonstrate the possibility of integrating off-the-shelf DB algorithm support with LLM agents through fine-grained adaptive interplay. Our framework, LEAD, enjoys both a regret guarantee and strong empirical performance.
@@ 4.1 Algorithmic Design of LEAD
In this section, we present the design intuitions of LEAD. We begin by discussing the desirable properties for an effective Algorithm-Enhanced LLM framework. Based on these considerations, we propose an agentic framework design LEAD, where we can incorporate any Explore-then-Exploit DB algorithm (Zoghi et al., 2014b) during inference. As an illustrative example, we use IF2 (Yue et al., 2012) to demonstrate how off-the-shelf algorithms can be integrated within LEAD and provide a detailed description.
Therefore, the Explore-Then-Exploit structure is particularly well-suited for LLMs (see Appendix B.1 for a detailed illustration). By selecting an Explore-Then-Exploit DB algorithm for LEAD, we address [Q1]. As an example, we use IF2 Yue et al. (2012) as a base algorithm to illustrate the theoretical guarantee and empirical performance. This approach can be applied similarly to other algorithms in the Explore-Then-Exploit family.
Algorithmic framework. The procedures of the LEAD are illustrated in Figure 4 and presented in Algorithm 1 (see more details in Appendix B.2). LEAD (IF2 base) maintains a confidence parameter $\delta$ and a threshold parameter $\epsilon$ that control the algorithm’s confidence of matches between arms. The key components of LEAD (IF2 base) are:
  • Phase 1 (LLM Phase): Utilization of LLM recommended arms. The agentic framework maintains a set of candidate arms $B$. Given two arms suggested by an LLM agent, the framework begins with finding a winner between them, denoted by $b_{\mathrm{LLM}}$. The winning arm $b_{\mathrm{LLM}}$ is then matched with each remaining arm $b\in B$. This phase continues until $b_{\mathrm{LLM}}$ is defeated or all arms in $B$ have been matched. The variable $\mathsf{TrustLLM}$ controls the execution of the LLM phase, and is set to $\mathtt{False}$ when $b_{\mathrm{LLM}}$ is defeated by another arm, indicating the LLM’s suggestions are no longer trusted.
  • Phase 2 (DB Phase): Roll back to IF2. If $b_{\mathrm{LLM}}$ is defeated, the framework switches to implement one round of IF2 with an incumbent arm $b_{\text{{IF2}}}$ selected from an estimated preference matrix $\hat{P}$.
4.2 Theoretical Guarantees for LEAD
In this section, we begin by identifying the vulnerability of using standalone LLM agents for dueling bandits in Theorem 4.1. Then we provide the theoretical guarantees of LEAD in Theorem 4.2 and  4.3, demonstrating its efficacy and convergence.
Theorem 4.1 (Vulnerability).
The proof of Theorem 4.1 is provided in Appendix B.3.2. This vulnerability highlights the need for a more robust approach to use in-context LLM agents while offering theoretical guarantees under diverse prompting scenarios.
Regret Bounds. Following the algorithmic design of LEAD in Section 4.1, LEAD (IF2 base) inherits the theoretical guarantees of IF2 (see Appendix B.3.1), while nontrivially leveraging the benefits of LLMs’ exceptional weak regret performance for exploration across a range of instances. Specifically, LEAD (IF2 base) has the following theoretical guarantee:
Theorem 4.2 (Expected Regret).
$\mathbb{E}\left[\mathsf{WR}(\textsc{LEAD})\right]$ $\leq\min\left\{\widetilde{O}\left(T_{\mathrm{LLM}}+\frac{K\log K}{\epsilon_{1,2}}\right),\right.$
$\quad\quad\quad\quad\left.\widetilde{O}\left(\frac{K\log T}{\epsilon_{1,2}}\right)\right\}.$ (2)
where $\widetilde{O}(\cdot)$ hides poly-logarithmic factors of $T$.
The following theorem indicates that the additional term $({K\log K})/{\epsilon_{1,2}}$ in (4.2) is almost tight. Its proof is provided in Appendix B.3.2.
Theorem 4.3 (Converse).
Given any algorithm ALG for dueling bandits, there exists an LLM agent recommending arms such that if ALG satisfies $\mathbb{E}\left[\mathsf{WR}(\textsc{ALG})\right]\leq T_{\mathrm{LLM}}$ (here, $T_{\mathrm{LLM}}$ is defined in Theorem 4.2.), then it must hold that ${\mathbb{E}\left[\mathsf{SR}(\textsc{ALG})\right]\geq\mathbb{E}\left[\mathsf{WR}(\textsc{ALG})\right]=\Omega\left(T\right)}$.
4.3 Empirical Evaluation of LEAD
4.3.1 Efficacy Evaluation: Strong Regret and Weak Regret
Hyperparameters. In our implementation of LEAD (see Algorithm 1), there are two hyper-parameters: the threshold parameter $t$, which controls the maximum number of comparisons between arms, and the confidence parameter $\delta$, which determines the confidence level for pruning suboptimal arms. For the threshold parameter $t$, we considered values from the set $\{50,100,200\}$, and for the confidence parameter $\delta$, we explored values from $\{0.1,0.2,0.4\}$. After fine-tuning, we found that setting $t=50$ and $\delta=0.4$ provided the best performance in terms of cumulative regret.
4.3.2 Robustness Evaluation: Noisy and Adversarial Prompts
Figure 6: Duel selection trajectory comparison between GPT-4 Turbo and LEAD under different prompt settings (see Figures 7 and 8). (i) Top Row: With the original prompt, LEAD leverages GPT-4 Turbo’s exploration ability and guarantees convergence through the IF2 phase. (ii) Middle Row: With a noisy prompt (biased history), LEAD overcomes the limitation of standalone GPT-4 Turbo getting trapped in local optima by employing uniform comparisons in the IF2 phase. (iii) Bottom Row: With an adversarial prompt (reversed goal), LEAD maintains robust behavior despite the adversarial modification. Across all scenarios, LEAD demonstrates superior performance and robustness compared to standalone GPT-4 Turbo.
Recent studies (Loya et al., 2023; Krishnamurthy et al., ) have emphasized the importance of varying prompts to elicit the desired behavior from LLMs in decision-making tasks, highlighting the potential limitations of prompt quality. Results obtained from a single prompt template may lead to unreliable conclusions that cannot be generalized to real-world situations where optimal prompts are often unavailable. Thus, we evaluate the robustness of LEAD by employing two types of prompt perturbations (see Figure 8) along with the original prompt (see Figure 7). Across all scenarios, LEAD demonstrates superior performance and robustness compared to standalone GPT-4 Turbo.
Original prompt. Under the initial prompt, LEAD leverages the LLM’s ability to quickly identify the best arm through exploration. As shown in Figure 6 (Top Row), we observe that LEAD benefits from the LLM’s exploration ability by initializing with the best arm as the incumbent when entering the DB phase. Compared to GPT-4 Turbo, convergence to the Condorcet winner is guaranteed for LEAD with high probability.
Biased history. We inject an incorrect history into the prompt, where each non-optimal arm initially wins against the best arm 10 times, while keeping the underlying preference matrix unchanged. LLM agents are observed to get trapped in local optima, where LEAD overcomes this by employing uniform comparisons in the DB phase to escape such suboptimal exploration modes.
Reversed goal. When the prompt is adversarially modified from maximizing reward to minimizing, the LLM consistently recommends non-optimal arms after its exploration stage. Even with adversarial prompts, LEAD still achieves near-optimal cumulative strong regret. Since the LLM’s exploration capability is used within the bounded length of the Match Arms procedure, the impact of this adversarial attack is mitigated.
Figure 5 (right) presents the cumulative strong regret results comparing LEAD against standalone LLM agents and the IF2 algorithm across three prompt designs. It is worth noting that LEAD with $\delta=1/(TK^{2})$ (consistent with IF2) achieves near-optimal cumulative regret with low variance even with noisy and adversarial prompts, validating the regret guarantee in Theorem 4.2. LEAD and IF2 converge to the best arm within 2000 steps, while GPT-4 Turbo’s cumulative expected regret continues to increase, revealing the instability of standalone in-context LLM agents.
5 Conclusion
This paper evaluates LLMs as in-context decision-makers for context-free dueling bandits (DB) with a Condorcet Winner, offering the first systematic insights into their strengths and limitations. Our findings reveal that LLMs’ decision-making in DB, driven by linguistic priors, achieves exceptional weak regret performance across both transitive and intransitive environment instances in the short-term. However, LLMs lack the necessary criteria for convergence and long-term generalization to hard scenarios, leading to an optimality gap between LLMs and classic DB algorithms in terms of strong regret. To bridge this gap, we propose LEAD, an agentic flow framework that integrates off-the-shelf DB algorithms with LLM agents through fine-grained adaptive interplay. This framework provides theoretical guarantees and ensures robust performance even under noisy and adversarial prompts. Moving beyond the limitations of traditional numeric rewards, it sheds light on how language-based reasoning can be robustly generalized from words to actions.
6 Limitations
This work primarily focuses on the fundamental context-free Dueling Bandit (DB) setting, exploring the performance and generalization of Large Language Models (LLMs) within this relative decision-making framework. The scope of this work is limited to this specific setting, and does not address the complexities introduced by other types of DB environments. The following limitations are noteworthy: (i) The exploration of LLMs in conjunction with other ongoing regret-minimization algorithms is not considered here; (ii) The performance of LLMs under alternative winner definitions, such as Borda and Neumann winners is not explored; (iii) The paper does not examine LLMs’ behavior in other DB settings, such as contextual dueling bandits, multi-dueling bandits, and adversarial dueling bandits. Expanding the scope to solve these limitations is an important direction for future work.
7 Acknowledgment
We would like to thank all the anonymous reviewers for their helpful comments. Tongxin Li is the corresponding author. This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant No. 72301234; the Guangdong Basic and Applied Basic Research Foundation (No. 2025A1515011311); the PengCheng Peacock Supporting Scientific Research Fund; and the Guangdong Provincial Key Laboratory of Mathematical Foundations for Artificial Intelligence (No. 2023B1212010001).
Appendix
  1. A.
    Related Works
  2. B.
    Theoretical Part: Algorithm Design and Analysis of LEAD
  3. C.
    Experimental Part: Prompt Design and Supplementary Results
    • Appendix C.1 illustrates the implementation details of our experiments.
Appendix A Related Works
We provide the detailed related works as follows.
Dueling bandits. The problem of dueling bandits was initially introduced in (Yue et al., 2012). Various methods have been proposed to tackle the task since then. These methods can be broadly classified into two categories as Explore-Then-Exploit methods and Ongoing Regret Minimization methods according to (Zoghi et al., 2014b). Explore-Then-Exploit methods focus on identifying the best arm with high confidence before exploiting it, such as Interleaved Filter (IF(Yue et al., 2012) and Beat the Mean (BTM(Yue and Joachims, 2011), etc. In contrast, Ongoing Regret Minimization methods explicitly target the objective of minimizing cumulative regret, including Relative Upper Confidence Bound (RUCB(Zoghi et al., 2014a) and Self-Sparring (Sui et al., 2017), etc. Dueling bandit problem and preference feedback in general has a wide variety of applications, including recommendation systems (Yue et al., 2012), robotics (Tucker et al., 2020), and most recently, the training algorithm of large language models, such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022).
LLM agents for multi-armed bandits. Several recent works have explored evaluating the capabilities of LLMs in bandit problems. For example, (Baheri and Alm, 2023) proposed an approach to enhance contextual bandits by integrating LLMs as encoders. The LLMs’ ability to capture rich semantic and syntactic information from textual contexts is leveraged to provide the algorithm with a more informative representation of the context. The LLM-augmented algorithm transforms the raw context into a latent space vector using the LLM’s encoding capabilities. This encoded context is then used to guide the decision-making process. (Krishnamurthy et al., ) investigates whether LLMs can engage in exploration in simple MAB environments without additional training. They compared various prompt designs and found that GPT-4 with zero-shot chain-of-thought (CoT) reasoning and an externally summarized interaction history performed the best, while other configurations failed in exploration, either by never selecting the best arm after initial rounds or by selecting all arms nearly equally often. Different from the previous results, in this work we go beyond the settings of numeric rewards and investigate the capabilities of LLMs under preference feedback.
In-context LLMs for decision-making. Beyond bandit problems, LLM agents have demonstrated strong capabilities in complex reasoning across a wide range of in-context reinforcement learning and decision-making tasks (Laskin et al., ; Lee et al., 2024; Zhou et al., ; Yao et al., 2024). Various existing works aim to understand LLM agents’ capabilities for in-context decision-making, with notable examples including planning (Huang et al., 2023; Hao et al., ). Additionally, LLM agents have been shown to enhance embodied agents in robotic applications by providing advanced task planning abilities (Brohan et al., 2023) and reward designing (Ma et al., 2023), further enabling the development of lifelong learning agents (Wang et al., ). Besides these empirical successes, the authors of (Park et al., ) analyzed LLMs’ interactions in online learning and game theory settings through the lens of the regret metrics. They identified simple cases where LLMs fail to be no-regret. Another line of research incorporates LLMs into classic decision-making frameworks to create LLM-augmented online decision-makers. For instance, Liu et al. (Liu et al., ) utilized LLMs to enhance the components of warm starting, sampling candidates, and surrogate modeling in Bayesian optimization. Our work contributes to this broad area by integrating LLM agents with the classic Explore-then-Exploit DB algorithms to enhance the utilization of preference feedback.
Appendix B Algorithm Design and Analysis of LEAD
B.1 Algorithm Design Logic
@@
Explore-then-exploit algorithms as ideal candidates. Classic DB algorithms can be classified into two categories: Explore-Then-Exploit methods and Ongoing Regret Minimization methods (Zoghi et al., 2014b). Among these, Explore-Then-Exploit structure stands out as particularly well-suited for LLM augmentation:
  • Its symbolic representation of the algorithm’s logic enables clear integration of LLM suggestions at specific points without disrupting the overall structure and theoretical guarantees. In contrast, algorithms like Self-Sparring in (Sui et al., 2017) are less symbolic, making them less suitable for direct LLM augmentation.
B.2 Detailed Procedure Description
@@
Remark 1.
Remark 2.
@@
In Procedure 1 below, we describe the Match Arms procedure used in LEAD (see Algorithm 1 and Figure 4). It compares two given arms $a$ and $a^{\prime}$ with specified parameters $\delta$ and $\epsilon$.
@@
Similarly, the next Validate procedure is used in LEAD. It validates if an incumbent arm $b^{\prime}$ that wins previous matches in LEAD can still be trusted.
@@
We also reprise the IF2 procedure in (Yue et al., 2012) below to complement the presentation of LEAD.